If A not equal to B but A^2 = 5B-– 3, B^2 = 5A-– 3, then the
equation having the roots A/B and B/A is
(A) 3x^2 + 19x + 3 = 0
(B) 3x^2+ –19x + 3 = 0
(C) 3x^2 –+19x – 3 = 0
(D) x^2 –+16x + 1 = 0

Grade:10

1 Answers

Nishant Vora IIT Patna
8 years ago
A^2 – 5B+– 3= 0,
B^2 – 5A + 3= 0

So both A and B satisfies x^2 – 5x + 3= 0
So A and B are roots of above eqn
Now from give info just find sum of roots = A/B + B/A
and Product of roots = A/B * B/A = 1

Hence find the eqn